OpenStudy (anonymous):

Simplify using Trig I.D's 2cos^2sin/cos

OpenStudy (ybarrap):

$$ 2\cos^2(x) \cfrac{\sin (x)}{\cos( x)}=2\cos(x)\cancel{\cos(x)} \cfrac{\sin (x)}{\cancel{\cos( x)}}=2\cos(x)\sin (x)=\sin(2x) $$ Let me know if you have any questions.